# ../svg_images/ Quantum Architecture Search with Unsupervised Representation Learning

### v3

@@ 4.2 Quantum Architecture Search (QAS) Performance
Observation (3): In Table 2, we compare various QAS methods with our approach on the 4-qubit state preparation task, using a circuit space of 100,000 circuits and limiting the search to 1000 queries. GNNURL and GSQASURL represent predictor-based methods from He et al. (2023b) and He et al. (2023a), respectively, both employing our pre-trained model. QAS denotes the QAS approach with REINFORCE (BO) used in this work. The average results over 50 runs indicate that both the predictor-based methods and our approach are capable of identifying a significant number of high-performance circuits with fewer samples. However, predictor-based methods rely on labeled circuits to train predictors, introducing uncertainty as they may inadvertently filter out well-performing architectures along with poor ones. While a higher value filters out more low-performance circuits, increasing the proportion of good architectures in the filtered space, it also sacrifices many well-performing circuits, which can lead to improved Random Search performance but at the cost of excluding some optimal circuits. Despite these trade-offs, our method achieves comparable performance to predictor-based methods, demonstrating higher efficiency in terms of while requiring fewer circuit evaluations. In Appendix A.4, we present the best candidate circuits acquired by each of the three methods for every experiment.
@@ 5 Conclusion
Inspired by the Arch2vec method (Yan et al., 2020), we focus on exploring whether unsupervised architecture representation learning can enhance QAS. By decoupling unsupervised architecture representation learning from the QAS process, we successfully eliminate the need for a large number of labeled circuits. Additionally, our proposed quantum circuit encoding scheme addresses limitations in existing representations, improving search performance through more accurate and effective embeddings. Furthermore, our framework conducts QAS without relying on a predictor by directly applying search algorithms, such as REINFORCE and Bayesian Optimization (BO), to the latent representations. We have demonstrated the effectiveness of this approach through various experiments. In our framework, the success of QAS depends on the quality of unsupervised architecture representation learning and the selection of search algorithms. Thus, we recommend further investigation into architecture representation learning for QAS, as well as the development of more efficient search strategies within the latent representation space.
@@ Acknowledgments
The project of this workshop paper is supported with funds from the German Federal Ministry of Education and Research in the funding program Quantum Reinforcement Learning for industrial Applications (QLindA) - under project number 13N15644 and the Federal Ministry for Economic Affairs and Climate Action in the funding program Quantum-Classical Hybrid Optimization Algorithms for Logistics and Production Line Management (QCHALLenge) - under project number 01MQ22008B. The sole responsibility for the paper’s contents lies with the authors.
@@ A.2 Application Settings
In quantum information theory, fidelity (Liang et al., 2019) is an important metric to measure the similarity of two quantum states. By introducing fidelity as the performance index, we aim to maximize the similarity of the final state density operator with a certain desired target state. We first obtain the target state by randomly generating a corresponding circuit, and then with a limited number of sample circuits, we use the search methods to search candidate circuits that can achieve a fidelity higher than a certain threshold. During the search process, the fidelity can be directly used as a normalized reward function since its range is [0, 1]. Figure 5 shows the circuits used to generate the corresponding target states.
@@
The max-cut problem (Poljak & Rendl, 1995) consists of finding a decomposition of a weighted undirected graph into two parts (not necessarily equal size) such that the sum of the weights on the edges between the parts is maximum. Over these years, the max-cut problem can be efficiently solved with quantum algorithms such as QAOA (Villalba-Diez et al., 2021) and VQE (using eigenvalues). In our work, we address the problem by deriving the Hamiltonian of the graph and using VQE to solve it. We use a simple graph with the ground state energy for the 4-qubit experiment and a relatively complex graph with the ground state energy in the case of the 8-qubit experiment. Furthermore, we convert the energy into a normalized reward function integral to the search process. The visual representations of these graphs are presented below:
@@
In the field of QC, VQE (Peruzzo et al., 2014; Tilly et al., 2022) is a hybrid quantum algorithm for quantum chemistry, quantum simulations, and optimization problems. It is used to compute the ground state energy of a Hamiltonian based on the variational principle. For the 4- and 12-qubit quantum chemistry experiment, we use the Hamiltonian of the molecule and and its common approximate ground state energy and as the optimal energy. As for the 8-qubit experiment, we consider transverse field Ising model (TFIM) with the Hamiltonian as follows:

### v4

@@ 4.2 Quantum Architecture Search (QAS) Performance
Observation (3): In Table 2, we compare various QAS methods with our approach on the 4-qubit state preparation task, using a circuit space of 100,000 circuits and limiting the search to 1000 queries. GNNURL and GSQASURL represent predictor-based methods from [49] and [20], respectively, both employing our pre-trained model. QAS denotes the QAS approach with REINFORCE (BO) used in this work. The average results over 50 runs indicate that both the predictor-based methods and our approach are capable of identifying a significant number of high-performance circuits with fewer samples. However, predictor-based methods rely on labeled circuits to train predictors, introducing uncertainty as they may inadvertently filter out well-performing architectures along with poor ones. While a higher value filters out more low-performance circuits, increasing the proportion of good architectures in the filtered space, it also sacrifices many well-performing circuits, which can lead to improved Random Search performance but at the cost of excluding some optimal circuits. Despite these trade-offs, our method achieves comparable performance to predictor-based methods, demonstrating higher efficiency in terms of while requiring fewer circuit evaluations. In Appendix A.4, we present the best candidate circuits acquired by each of the three methods for every experiment.
@@
4.3 Evaluation on Real Quantum Hardware
To evaluate the deployability of the circuits discovered by our architecture search, we perform additional experiments using IBM’s real quantum device ibm_sherbrooke. For both the 4-node and 8-node MaxCut tasks, we selected the best-performing circuits and trained their parameters on an noise-free simulator. Once trained, we directly executed these circuits—without any further optimization—on the real device using the same parameters.
Observation.
As shown in Table 4, despite the presence of hardware noise and decoherence in the NISQ device, both MaxCut circuits retained their optimal output performance when transferred from simulator to real hardware. The circuits achieved a 100% probability of measuring the optimal bitstring in 10,000 repeated shots, identical to the simulator outcome. These results validate that the discovered quantum architectures not only perform well in idealized environments but also translate reliably to real-world quantum processors without requiring parameter re-tuning.
@@ 5 Conclusion
Metric | 4-node MaxCut | 8-node MaxCut | ||
Simulator | Real Device | Simulator | Real Device | |
Probability of Optimal Bitstring | 100.0% | 100.0% | 100.0% | 100.0% |
Average Cut Value | 4 | 4 | 12 | 12 |
Number of Shots | 10,000 | 10,000 | 10,000 | 10,000 |
Noise Model | Noise-free | Physical (NISQ) | Noise-free | Physical (NISQ) |
Device | Simulator | ibm_sherbrooke | Simulator | ibm_sherbrooke |
In this work, we focus on exploring whether unsupervised architecture representation learning can enhance QAS. By decoupling unsupervised architecture representation learning from the QAS process, we successfully eliminate the need for a large number of labeled circuits. Additionally, our proposed quantum circuit encoding scheme addresses limitations in existing representations, improving search performance through more accurate and effective embeddings. Furthermore, our framework conducts QAS without relying on a predictor by directly applying search algorithms, such as REINFORCE and Bayesian Optimization (BO), to the latent representations. We have demonstrated the effectiveness of this approach through various experiments on simulator and real quantum hardware. In our framework, the success of QAS depends on the quality of unsupervised architecture representation learning and the selection of search algorithms. Thus, we recommend further investigation into architecture representation learning for QAS, as well as the development of more efficient search strategies within the latent representation space.
@@ Acknowledgments
The project of this workshop paper is supported with funds from the German Federal Ministry of Education and Research in the funding program Quantum Reinforcement Learning for industrial Applications (QLindA) - under project number 13N15644 and the Federal Ministry for Economic Affairs and Climate Action in the funding program Quantum-Classical Hybrid Optimization Algorithms for Logistics and Production Line Management (QCHALLenge) - under project number 01MQ22008B. The sole responsibility for the paper’s contents lies with the authors. Portions of this manuscript were generated with the assistance of OpenAI’s ChatGPT-4 model for drafting and language refinement. The authors take full responsibility for the content.
@@ A.2 Application Settings
In quantum information theory, fidelity [21] is an important metric to measure the similarity of two quantum states. By introducing fidelity as the performance index, we aim to maximize the similarity of the final state density operator with a certain desired target state. We first obtain the target state by randomly generating a corresponding circuit, and then with a limited number of sample circuits, we use the search methods to search candidate circuits that can achieve a fidelity higher than a certain threshold. During the search process, the fidelity can be directly used as a normalized reward function since its range is [0, 1]. Figure 5 shows the circuits used to generate the corresponding target states.
@@
The max-cut problem [22] consists of finding a decomposition of a weighted undirected graph into two parts (not necessarily equal size) such that the sum of the weights on the edges between the parts is maximum. Over these years, the max-cut problem can be efficiently solved with quantum algorithms such as QAOA [58] and VQE (using eigenvalues). In our work, we address the problem by deriving the Hamiltonian of the graph and using VQE to solve it. We use a simple graph with the ground state energy for the 4-qubit experiment and a relatively complex graph with the ground state energy in the case of the 8-qubit experiment. Furthermore, we convert the energy into a normalized reward function integral to the search process. The visual representations of these graphs are presented below:
@@
In the field of QC, VQE [13, 12] is a hybrid quantum algorithm for quantum chemistry, quantum simulations, and optimization problems. It is used to compute the ground state energy of a Hamiltonian based on the variational principle. For the 4- and 12-qubit quantum chemistry experiment, we use the Hamiltonian of the molecule and and its common approximate ground state energy and as the optimal energy. As for the 8-qubit experiment, we consider transverse field Ising model (TFIM) with the Hamiltonian as follows: